OpenStudy (anonymous):

B = 703 * w/h (squared) Note: w/h is a fraction where h is squared Part A: Solve for W

OpenStudy (anonymous):

okay, so the whole thing is divided by h^2

sam (.sam.):

\[B=\frac{W}{h^2}\] Multiply both sides by h^2 \[Bh^2=W\] Substitute B and h into equation \[W=(21.45)(64)^2\] Solve

OpenStudy (anonymous):

Thank you Sam for helping me with Part B. I still need help with Part A which is: \[B = 703 * \left(\begin{matrix}w \\ h ^{2}\end{matrix}\right)\] , solve for W

sam (.sam.):

Part A B = 703 * w/h multiply both sides by h Bh=703w divide both sides by 703 \[w=\frac{Bh}{703}\]

OpenStudy (anonymous):

the h is squared though. So, I would divide both sides by h squared instead, right?

sam (.sam.):

yes

sam (.sam.):

Part A \[B = 703 \times \frac{w}{h^2}\] multiply both sides by \(h^2\) \[Bh^2=703w \] divide both sides by 703 \[w=\frac{Bh^2}{703} \]

sam (.sam.):

The 703 is given or you put it?

OpenStudy (anonymous):

It is given as a 2 part question. I will write it how it looks ...... Question 16: "The formula to compute a person's body mass index is \[B = 703 * \left(\begin{matrix}w \\ h ^{2}\end{matrix}\right)\] B represents the body mass index. W is the person's weight in pounds, and h represents the person's height in inches. Part A: "Solve for W" (which you already helped me understand) Part B: "What is the weight (to the nearest pound) of a person who is 64 inches tall and has a body mass index of 21.45?"

sam (.sam.):

For part B \[B=703 \times \frac{w}{h^2}\] multiply both sides by h^2 \[Bh^2=703w\] divide both sides by 703 \[w=\frac{Bh^2}{703}\] -------------------------------------------- B=21.45 , h= 64 \[w=\frac{(21.45)(64)^2}{703}\]

sam (.sam.):

w=125

sam (.sam.):

I'll be back in 2 hours, the answer for part B is my previous post

OpenStudy (anonymous):

Thank you so much. I just figured it out with the calculator. By writing it down by hand step by step how you showed me. I now can understand it. Thanks. Are you available to check my work on another problem?